# Superintegrability of the monomial Uglov matrix model

### v1

@@ 1 Introduction
This paper is devoted to the construction of new exactly solvable matrix, or rather, eigenvalue models. They are not just interesting as a new example, but also for the kinds of structures that they are related to. For instance, to construct these new examples, we apply our intuition about hidden symmetries responsible for superintegrability. Superintegrability seems to be tied to the theory of characters and in general symmetric functions naturally labeled by partitions. In fact it was discovered recently that the combinatorial nature of the exact solutions for the averages of these symmetric functions has an algebraic structure behind it. In particular, an important role is played by algebras which belong to the family of so-called BPS-algebras [22, 23, 24]. The relevance of representations of the $\mathcal{W}_{1+\infty}$ algebra, the Affine Yangian $Y\left(\hat{\mathfrak{gl}}_{1}\right)$ [25, 26, 27, 28] where observed extensively [29, 30]. Superintegrability appears as a consequence of special representations of these algebras with states indexes by partitions. On the one hand, questions arising in matrix models can be used to establish further understanding of these algebras and motivate questions. This path was taken, for example, in [31, 32] where certain commutative subalgebras of $Y\left(\hat{\mathfrak{gl}}_{1}\right)$ were identified and were related to novel integrable systems. On the other hand as we demonstrate here one can borrow ideas from the relevant representation theory to extend the notion of superintegrability to new examples.
@@
We would like to state the main result here and refer to bulk of the paper for the explanation of notations. It is well known that an explicit formula exists for the monomial matrix models. Namely, in that case for Schur polynomials one has [36]:
@@
We call this eigenvalue model the monomial Uglov model, following [37] where the Uglov matrix model has been introduced. One can see that the measure $\mu(x_{i},x_{j})$ in this case is more complicated. The basis for superintegrability is now given by the so-called Uglov polynomials, which are are a deformation of Schur polynomials, similar to Jack polynomials. They reduce, together with the measure, and, hence, the whole formula to the undeformed case at $\beta\rightarrow 1$. In the paper we also propose a second version of this formula for a slightly different measure. Let us point out a few interesting relations that this model and it’s components have to other subjects.
@@
  • The key component is the Uglov limit. The Uglov polynomials where originally introduced by D.Uglov in [38] as a special limit of Macdonald polynomials, with $(q,t)$ around a primary root of unity point. There they where shown to be in one-to-one correspondence with the eigenfunctions of the $\mathfrak{gl}_{s}$ spin-Calogero model. This is interesting on it’s own, as the relevance of the ordinary trigonometric Calogero system for $\beta$ deformed matrix models was observed in various scenarios [30]. They also recently appeared in various other contexts related to conformal field theory and representation theory of the super-Virasoro algebra [39, 40]
@@
  • As it was shown recently the statement above could be further uplifted. In case of Jack polynomials they can be shown to not only be eigenfunctions of the Calogero Hamiltonian, but actually states in the Fock representation of the full $Y\left(\mathfrak{\hat{gl}}_{1}\right)$ affine Yangian. Uglov polynomials, on the other supposedly play the same role in higher rank affine Yangians $Y\left(\mathfrak{\hat{gl}}_{s}\right)$ (see [41] where rank two $s=2$ was discussed, and also [42]).
@@
  • The measure of the Uglov matrix model was first introduced in [43]. It was motivated by considering the matrix models representation of the Nekrasov partition function on orbifolds $\mathbb{C}^{2}/\Gamma_{s,v}$.  [^3] We use a slightly different notation compared to [43] as in our case we preferred to keep $s$ for the degree of the potential.. As we will also stress later our results only apply to the cases, which correspond to $\mathbb{C}^{2}/\mathbb{Z}_{s}$ and ×(/s)$\mathbb{C}\times\left(\mathbb{C}/\mathbb{Z}_{s}\right)$. The potential for the instanton matrix model is more complicated to study in contexts of superintegrability. Various discussions about the relevance of the root of unity limit where already presented [36]. We hope our works serves as a further clarification of the subject. The relevance of the root of unity limit for gauge theories on orbifolds has be demonstrated in numerous works [44, 45, 46]
@@
The way that one can come up with this formula is interesting on itself. Let us briefly summarize the logic here. First, one has to somehow come up with the correct measure. Even then, as we will explain in the paper it is not possible to extend methods like [13, 29] to just derive the solution of the Uglov matrix model straightforwardly from Virasoro constraints. On the other hand, interpolating at different integer $N$ gets quite complicated quite fast, hence without knowing the correct basis of Uglov polynomials it does not seem plausible to guess it. Hence this is a very illuminating test case for using hidden symmetries instead. The key is guessing that the root of unity limit:
@@
is relevant to the non-Gaussian model. To do this one has to turn to techniques developed in [30]. Using techniques developed there one can reverse engineer the $W$-operator, that generated the undeformed monomial model. That allows to identify it as a root of unity limit of certain element of the quantum toroidal DIM algebra [47, 48, 20], with the $(q,t)$-dependent Macdonald operator being it’s building block. On the algebraic side, the relevance of the root of unity limit becomes clear. Even without explicitly studying the form of the deformed $W$-operator we then conjecture the superintegrability of the Uglov model. Let us mention here, that the root of unity limit of quantum algebras is interesting on its own [49, 50] and it’s role in the matrix model should be studied further. The described logic can be summarized by the following picture:
@@
A key difficulty here is that one cannot prove this statement by most methods, that work in the undeformed case. In particular in [36] the proof utilized the technique of orthogonal polynomials, which breaks down after $\beta$-deformation, just as all the methods related to free fermions and KP/Toda integrability. Instead below we provide evidence for it’s validity that we were able to find.
@@
In section 3 we introduce the main players of our construction: the Uglov matrix model and the Uglov polynomials. We discus the peculiarities of the root of unity limit. In fact, the limit produces a whole bunch of measures, indexed by an additional parameter $v$. Whereas taking the same limit in Macdonald polynomials seems to be problematic except for two special cases $v=0$ and $v=1$.
@@ 2 Superintegrability and the monomial matrix model
Here we have used a notation for Schur functions evaluated at special loci of their variables, which we explain in detail in Appendix A. The rightmost expression represents the $N$ dependence of the expectation value as a product of simple factors over the boxes of the Young diagram $R$. The value of $(j-i)$ is called the content of a box in the Young diagram. For that reason we will sometimes refer to such parts of formulas as the content product. Since superintegrability of matrix models is not yet a completely formalised property one can think of both the middle and the rightmost parts of the formula as it’s manifestation.
@@
where $\Delta(x)=\prod\limits_{i<j}(x_{i}-x_{j})$ is the Vandermonde determinant. The deformation is introduced by switching a the parameter in the power of the Vandermonde determinant:
@@
As we can see both manifestations of superintegrability get deformed. The content product now also depends on $\beta$. This is now a more peculiar statement, since at generic $\beta$ we cannot use Wick’s theorem effectively. Moreover lots of other techniques such as unitary Harish-Chandra–Itzykson–Zuber integration, orthogonal polynomials and others, related to free fermion representation and KP/Toda integrability of matrix models break down. The only straightforward way to deduce this formula, would be by interpolating at multiple integer values of $\beta$. Instead, a proof of these formulas required to look at algebraic structures behind superintegrability, which in this case appears to be the affine Yangian $Y(\hat{\mathfrak{gl}}_{1})$ [30, 29, 52]
@@
Another feature of the result (15) is a lack of representation of the content product in terms of special values of Schur functions. As announced in the introduction we will propose such formula below. Finally, to see how it reduces to the Gaussian case one has to simply put $s=2$ and $a=1$. In this case for each box the entry is either even or odd, hence either only the first or only the second factor is non-trivial. In that manner we obtain the product (8).
@@ 3 The Uglov limit and the Uglov matrix model
As we will explicitly demonstrate below, the naive $\beta$ deformation with substitution of the Vandermonde and Schur function for Jack polynomials does not work. Instead we should introduce a new measure and a new set of symmetric functions. In this section we describe both in detail and explain the motivation later. In both cases they can be identified as a certain limit of the respective objects depending on two parameters $(q,t)$ related to the quantum toroidal algebra. In this limit both parameters go to a root of unity with a certain scaling:
@@ 3.1 The Uglov Matrix model and root of unity limit of the $(q,t)$-matrix model
For brevity we will call this measure an Uglov determinant by analogy with the Vandermonde determinant. As announced above we will deal with two cases $v=0,v=1$ when:
@@
The origin of this measure as suggested by the limit (17) should be related the $\left(q,t\right)$-deformed matrix model. We will not deal with the whole $(q,t)$-models right now, but only with the measure part. Here we will briefly recall the derivation, as it was already done in [37]. The $(q,t)$-models are defined by substituting the Vandermonde determinant by:
@@ 3.2 Uglov polynomials
The key feature of Uglov limit, which was crucial for [38] is that it is consistent and well defined for Macdonald polynomials.
Denote the Macdonald polynomials as $\text{Mac}_{R}(p_{k}|q,t)$, depending on the integer partition $R$. Now we want to define the so-called Uglov polynomials as the limit (17) of Macdonald polynomials. Here we are forced to consider $v=0$ or $v=1$ as only in these cases the limit seems to be well defined. It is interesting to study whether extensions to other values of $v$ are possible. Hence we define:
@@
wherever the value of $\beta$ is not important. There are several comments to make here. First, at $v=1$ we could have set $q=\omega_{r}^{l}u,q=\omega_{r}^{l}s^{\beta}$. In this case the resulting polynomial does not depend on $l$. For $r=1$ these polynomials are nothing but Jack polynomials:
@@
The listed Uglov polynomials for $v=1$ don’t look that different from each other and Jack polynomials. It becomes more visible for bigger diagrams, and, what is more important is that properties of these functions are very different. We list some more examples in Appendix B. Some properties of Uglov polynomials as symmetric functions are collected in the Appendix in [41] and some will be revealed below.
@@
After looking at this table and a number of other examples one can notice that for diagrams which are divisible by $r$, i.e. such that all their row lengths are divisible by $r$, the respective polynomials only depend powers sums $p_{kr}$. Moreover one can go even further and conjecture the following formula:
@@
where $r\cdot R$ is the partition with all row lengths multiplied by $r$, i.e. $[rR_{1},rR_{2},\ldots]$. We will see below how this statement interacts nicely with the Cauchy formula and later superintegrability. For example one can take $R=[2,1]$ and $r=2$:
@@
An important formula in the theory of symmetric functions that we will use below is the so called Cauchy formula. For Schur polynomials it has the standard form [57]:
@@
Let us go into detail here to demonstrate once again how the dependence on $\beta\mod r$ arises when taking the limit. In particular if we denote by $\eta_{k}$ the coefficient in front of $\frac{p_{k}q_{k}}{k}$ in the l.h.s. of the Cauchy formula we have at $v=1$:
@@
where we parameterized $u$ from (17) as $u=e^{h}$, where $h\to 0$. Expanding explicitly one has:
@@
Hence the generating function depends only on powers sums with indices that are divisible by $r$. As we see this agrees nicely with the fact that in this case the symmetric polynomials $U_{R}^{(r,0)}$ with diagrams that are divisible by $r$ only depend on $p_{kr}$.
@@ 4.1 The main statement
We would like to briefly discuss the Gaussian case separately as it is the simplest example, and also involves an interesting feature. Notice that at $s=r=2$ the result for $v=1$ (49) expresses superintegrability of Uglov polynomials $U^{(2,1)}_{R}$ in the model given by:
@@
However, we know that at the same time the Gaussian weight can be integrated with the standard $\beta$-deformed Vandermonde and with a Jack polynomial insertion. In terms of the Uglov model this corresponds to taking $r=1$ and $s=2$. This means that there are two exactly solvable eigenvalue models available with the Gaussian potential. Consider for example the two expression for $R=[2]$, where formulas are very similar, and take $N=1\mod 2$:
@@
Apart from the Gaussian case, we could not find any superintegrability property when $r\neq s$, at least with Uglov polynomials as symmetric functions. We assume the following simple check. If the model is superintegrable we expect the averages to look like $\Big{\langle}U_{R}\Big{\rangle}={P(N,\beta)\over Q(\beta)}$, where $P$ is polynomial of $N$ and $\beta$, $Q$ is a polynomial only of $\beta$ and moreover we expect both to be factorized in $N$ and $\beta$. Moreover suppose the Uglov polynomials where not the right basis in this case. Then if there exists some other basis with superintegrable averages of the described form, then one could generically expect any average, being expended into a superintegrable basis to be of similar form. With that in mind we evaluate expectation values of various types, say, of $\Big{\langle}U_{R}^{(s,1)}\Big{\rangle}\,,\,\Big{\langle}U^{(r,1)}_{R}\Big{\rangle}$ of just powers sums $\Big{\langle}p_{\Delta}\Big{\rangle}$. In all cases we find that, for example, the denominator increases much faster then expected and contains large prime numbers as factors, which clearly cannot appear from simple factorized polynomials in $\beta$.
Still, these are indirect arguments, that illustrate that naive attempts to extend superintegrability beyond $r=s$ where unsuccessful.

### v2

@@ 1 Introduction
This paper is devoted to the construction of new exactly solvable matrix, or rather, eigenvalue models. They are not just interesting as a new example, but also for the kinds of structures that they are related to. For instance, to construct these new examples, we apply our intuition about hidden symmetries responsible for superintegrability. Superintegrability seems to be tied to the theory of characters and in general symmetric functions naturally labelled by partitions. In fact, it was discovered recently that the combinatorial nature of the exact solutions for the averages of these symmetric functions has an algebraic structure behind it. In particular, an important role is played by algebras which belong to the family of so-called BPS-algebras [22, 23, 24]. The relevance of representations of the $\mathcal{W}_{1+\infty}$ algebra, the Affine Yangian $Y\left(\hat{\mathfrak{gl}}_{1}\right)$ [25, 26, 27, 28] were observed extensively [29, 30]. Superintegrability appears as a consequence of special representations of these algebras with states indexes by partitions. On the one hand, questions arising in matrix models can be used to establish further understanding of these algebras and motivate questions. This path was taken, for example, in [31, 32] where certain commutative subalgebras of $Y\left(\hat{\mathfrak{gl}}_{1}\right)$ were identified and were related to novel integrable systems. On the other hand as we demonstrate here one can borrow ideas from the relevant representation theory to extend the notion of superintegrability to new examples.
@@
We would like to state the main result here and refer to the bulk of the paper for the explanation of notations. It is well known that an explicit formula exists for the monomial matrix models. Namely, in that case for Schur polynomials one has [36]:
@@
We call this eigenvalue model the monomial Uglov model, following [37] where the Uglov matrix model has been introduced. One can see that the measure $\mu(x_{i},x_{j})$ in this case is more complicated. The basis for superintegrability is now given by the so-called Uglov polynomials, which are a deformation of Schur polynomials, similar to Jack polynomials. They reduce, together with the measure, and, hence, the whole formula to the undeformed case at $\beta\rightarrow 1$. In the paper we also propose a second version of this formula for a slightly different measure. Let us point out a few interesting relations that this model and its components have to other subjects.
@@
  • The key component is the Uglov limit. The Uglov polynomials were originally introduced by D.Uglov in [38] as a special limit of Macdonald polynomials, with $(q,t)$ around a primary root of unity point. There they were shown to be in one-to-one correspondence with the eigenfunctions of the $\mathfrak{gl}_{s}$ spin-Calogero model. This is interesting on its own, as the relevance of the ordinary trigonometric Calogero system for $\beta$ deformed matrix models was observed in various scenarios [30]. They also recently appeared in various other contexts related to conformal field theory and representation theory of the super-Virasoro algebra [39, 40]
@@
  • As it was shown recently the statement above could be further uplifted. In the case of Jack polynomials they can be shown to not only be eigenfunctions of the Calogero Hamiltonian, but actually states in the Fock representation of the full $Y\left(\mathfrak{\hat{gl}}_{1}\right)$ affine Yangian. Uglov polynomials, on the other supposedly play the same role in higher rank affine Yangians $Y\left(\mathfrak{\hat{gl}}_{s}\right)$ (see [41] where rank two $s=2$ was discussed, and also [42]).
@@
  • The measure of the Uglov matrix model was first introduced in [43]. It was motivated by considering the matrix models representation of the Nekrasov partition function on orbifolds $\mathbb{C}^{2}/\Gamma_{s,v}$.  [^3] We use a slightly different notation compared to [43] as in our case we preferred to keep $s$ for the degree of the potential.. As we will also stress later, our results only apply to the cases, which correspond to $\mathbb{C}^{2}/\mathbb{Z}_{s}$ and ×(/s)$\mathbb{C}\times\left(\mathbb{C}/\mathbb{Z}_{s}\right)$. The potential for the instanton matrix model is more complicated to study in contexts of superintegrability. Various discussions about the relevance of the root of unity limit were already presented [36]. We hope our works serves as a further clarification of the subject. The relevance of the root of unity limit for gauge theories on orbifolds has been demonstrated in numerous works [44, 45, 46]
@@
The way that one can come up with this formula is interesting in itself. Let us briefly summarize the logic here. First, one has to somehow come up with the correct measure. Even then, as we will explain in the paper it is not possible to extend methods like [13, 29] to just derive the solution of the Uglov matrix model straightforwardly from Virasoro constraints. On the other hand, interpolating at different integer $N$ gets quite complicated quite fast, hence without knowing the correct basis of Uglov polynomials it does not seem plausible to guess it. Hence this is a very illuminating test case for using hidden symmetries instead. The key is guessing that the root of unity limit:
@@
is relevant to the non-Gaussian model. To do this one has to turn to techniques developed in [30]. Using techniques developed there one can reverse engineer the $W$-operator, that generated the undeformed monomial model. That allows to identify it as a root of unity limit of certain element of the quantum toroidal DIM algebra [47, 48, 20], with the $(q,t)$-dependent Macdonald operator being its building block. On the algebraic side, the relevance of the root of unity limit becomes clear. Even without explicitly studying the form of the deformed $W$-operator we then conjecture the superintegrability of the Uglov model. Let us mention here, that the root of unity limit of quantum algebras is interesting on its own [49, 50] and its role in the matrix model should be studied further. The described logic can be summarized by the following picture:
@@
A key difficulty here is that one cannot prove this statement by most methods, that work in the undeformed case. In particular in [36] the proof utilized the technique of orthogonal polynomials, which breaks down after $\beta$-deformation, just as all the methods related to free fermions and KP/Toda integrability. Instead, below we provide evidence for its validity that we were able to find.
@@
In section 3 we introduce the main players of our construction: the Uglov matrix model and the Uglov polynomials. We discuss the peculiarities of the root of unity limit. In fact, the limit produces a bunch of measures, indexed by an additional parameter $v$. Whereas taking the same limit in Macdonald polynomials seems to be problematic except for two special cases $v=0$ and $v=1$.
@@ 2 Superintegrability and the monomial matrix model
Here we have used a notation for Schur functions evaluated at special loci of their variables, which we explain in detail in Appendix A. The rightmost expression represents the $N$ dependence of the expectation value as a product of simple factors over the boxes of the Young diagram $R$. The value of $(j-i)$ is called the content of a box in the Young diagram. For that reason we will sometimes refer to such parts of formulas as the content product. Since superintegrability of matrix models is not yet a completely formalised property one can think of both the middle and the rightmost parts of the formula as its manifestation.
@@
where $\Delta(x)=\prod\limits_{i<j}(x_{i}-x_{j})$ is the Vandermonde determinant. The deformation is introduced by switching a parameter in the power of the Vandermonde determinant:
@@
As we can see, both manifestations of superintegrability get deformed. The content product now also depends on $\beta$. This is now a more peculiar statement, since at generic $\beta$ we cannot use Wick’s theorem effectively. Moreover, lots of other techniques such as unitary Harish-Chandra–Itzykson–Zuber integration, orthogonal polynomials and others, related to free fermion representation and KP/Toda integrability of matrix models break down. The only straightforward way to deduce this formula, would be by interpolating at multiple integer values of $\beta$. Instead, a proof of these formulas required to look at algebraic structures behind superintegrability, which in this case appears to be the affine Yangian $Y(\hat{\mathfrak{gl}}_{1})$ [30, 29, 52]
@@
Another feature of the result (15) is a lack of representation of the content product in terms of special values of Schur functions. As announced in the introduction, we will propose such formula below. Finally, to see how it reduces to the Gaussian case, one has to simply put $s=2$ and $a=1$. In this case, for each box the entry is either even or odd, hence either only the first or only the second factor is non-trivial. In that manner, we obtain the product (8).
@@ 3 The Uglov limit and the Uglov matrix model
As we will explicitly demonstrate below, the naive $\beta$ deformation with substitution of the Vandermonde and Schur function for Jack polynomials does not work. Instead, we should introduce a new measure and a new set of symmetric functions. In this section we describe both in detail and explain the motivation later. In both cases they can be identified as a certain limit of the respective objects depending on two parameters $(q,t)$ related to the quantum toroidal algebra. In this limit both parameters go to a root of unity with a certain scaling:
@@ 3.1 The Uglov Matrix model and root of unity limit of the $(q,t)$-matrix model
For brevity we will call this measure an Uglov determinant by analogy with the Vandermonde determinant. As announced above, we will deal with two cases $v=0,v=1$ when:
@@
The origin of this measure as suggested by the limit (17) should be related to the $\left(q,t\right)$-deformed matrix model. We will not deal with the whole $(q,t)$-models right now, but only with the measure part. Here we will briefly recall the derivation, as it was already done in [37]. The $(q,t)$-models are defined by substituting the Vandermonde determinant by:
@@ 3.2 Uglov polynomials
The key feature of Uglov limit, which was crucial for [38] is that it is consistent and well-defined for Macdonald polynomials.
Denote the Macdonald polynomials as $\text{Mac}_{R}(p_{k}|q,t)$, depending on the integer partition $R$. Now we want to define the so-called Uglov polynomials as the limit (17) of Macdonald polynomials. Here we are forced to consider $v=0$ or $v=1$ as only in these cases the limit seems to be well-defined. It is interesting to study whether extensions to other values of $v$ are possible. Hence, we define:
@@
wherever the value of $\beta$ is not important. There are several comments to make here. First, at $v=1$ we could have set $q=\omega_{r}^{l}u,q=\omega_{r}^{l}s^{\beta}$. In this case, the resulting polynomial does not depend on $l$. For $r=1$ these polynomials are nothing but Jack polynomials:
@@
The listed Uglov polynomials for $v=1$ don’t look that different from each other and Jack polynomials. It becomes more visible for bigger diagrams, and, what is more important, is that properties of these functions are very different. We list some more examples in Appendix B. Some properties of Uglov polynomials as symmetric functions are collected in the Appendix in [41] and some will be revealed below.
@@
After looking at this table and a number of other examples one can notice that for diagrams which are divisible by $r$, i.e. such that all their row lengths are divisible by $r$, the respective polynomials only depend on power sums $p_{kr}$. Moreover one can go even further and conjecture the following formula:
@@
where $r\cdot R$ is the partition with all row lengths multiplied by $r$, i.e. $[rR_{1},rR_{2},\ldots]$. We will see below how this statement interacts nicely with the Cauchy formula and later superintegrability. For example, one can take $R=[2,1]$ and $r=2$:
@@
An important formula in the theory of symmetric functions that we will use below is the so-called Cauchy formula. For Schur polynomials, it has the standard form [57]:
@@
Let us go into detail here to demonstrate once again how the dependence on $\beta\mod r$ arises when taking the limit. In particular, if we denote by $\eta_{k}$ the coefficient in front of $\frac{p_{k}q_{k}}{k}$ in the l.h.s. of the Cauchy formula we have at $v=1$:
@@
where we parameterized $u$ from (17) as $u=e^{h}$, where $h\to 0$. Expanding explicitly, one has:
@@
Hence the generating function depends only on powers sums with indices that are divisible by $r$. As we see, this agrees nicely with the fact that in this case the symmetric polynomials $U_{R}^{(r,0)}$ with diagrams that are divisible by $r$ only depend on $p_{kr}$.
@@ 4.1 The main statement
We would like to briefly discuss the Gaussian case separately, as it is the simplest example, and also involves an interesting feature. Notice that at $s=r=2$ the result for $v=1$ (49) expresses superintegrability of Uglov polynomials $U^{(2,1)}_{R}$ in the model given by:
@@
However, we know that at the same time the Gaussian weight can be integrated with the standard $\beta$-deformed Vandermonde and with a Jack polynomial insertion. In terms of the Uglov model, this corresponds to taking $r=1$ and $s=2$. This means that there are two exactly solvable eigenvalue models available with the Gaussian potential. Consider for example the two expression for $R=[2]$, where formulas are very similar, and take $N=1\mod 2$:
@@
Apart from the Gaussian case, we could not find any superintegrability property when $r\neq s$, at least with Uglov polynomials as symmetric functions. We assume the following simple check. If the model is superintegrable we expect the averages to look like $\Big{\langle}U_{R}\Big{\rangle}={P(N,\beta)\over Q(\beta)}$, where $P$ is a polynomial of $N$ and $\beta$, $Q$ is a polynomial only of $\beta$ and moreover we expect both to be factorized in $N$ and $\beta$. Moreover, suppose the Uglov polynomials were not the right basis in this case. Then if there exists some other basis with superintegrable averages of the described form, then one could generically expect any average, being expended into a superintegrable basis to be of similar form. With that in mind, we evaluate expectation values of various types, say, of $\Big{\langle}U_{R}^{(s,1)}\Big{\rangle}\,,\,\Big{\langle}U^{(r,1)}_{R}\Big{\rangle}$ of just powers sums $\Big{\langle}p_{\Delta}\Big{\rangle}$. In all cases we find that, for example, the denominator increases much faster than expected and contains large prime numbers as factors, which clearly cannot appear from simple factorized polynomials in $\beta$.
Still, these are indirect arguments, that illustrate that naive attempts to extend superintegrability beyond $r=s$ were unsuccessful.